# On the existence of weighted-cscK metrics

### v2

@@ 1. Introduction
In this paper, we apply the variational approach and continuity method to study the weighted-cscK (extremal) metrics, the study of which is initiated in papers [Lah19, Lah23, AJL23]. The weighted-cscK (extremal) metric is a natural generalization of the cscK (extremal) metric. The (generalized) Yau–Tian–Donaldson conjecture predicts an equivalence between the existence of cscK metric(s) and the ${\mathbb{G}}$-uniform K-stability condition. By the work of many authors (see an incomplete list: [Tia97, DR17, BDL20, CC21a, CC21b, Li22, BJ23]), much progress has been made, and it is known that the existence of cscK metric(s) is equivalent to the ${\mathbb{G}}$-coercivity of the Mabuchi functional, and can be implied by the ${\mathbb{G}}$-uniformly K-stability for models.
@@
$\mathrm{tr}_{\mathrm{v},\phi}(\eta):=\mathrm{v}(m_{\phi})\mathrm{tr}_{\phi}\eta+\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\eta}\rangle,$
where $\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\eta}\rangle$ is uniformly bounded independent of $\phi$. For the detail definition of $m_{\eta}$ see Subsection 2.A.
@@
To show (2) implies (1), we consider the following weighted version of the continuity path of Chen [Che18] for weighted extremal metrics,
@@
where $\theta$ is a $K$-invariant smooth positive $(1,1)$-form in the same class of $\omega$.
@@
The idea of the continuity path (1.3) can be seen from the variational picture. Since $-\mathrm{Ric}(\omega)$ may not be positive, replacing a small part of the term $-{\mathbb{E}}_{\mathrm{v}}^{\mathrm{Ric}}$ in the weighted Mabuchi functional by ${\mathbb{E}}_{\mathrm{v}}^{\theta}$ for some positive (1,1)-form $\theta$ will make the weighted Mabuchi functional more coercive. Intuitively, this makes (1.3) always solvable for $t\in[0,1)$.
@@
$\qquad\qquad\qquad\qquad\qquad\qquad-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})}.$ (1.4b)
@@
We prove a uniform $C^{0}$-estimate by modifying Guo-Phong’s argument in [GP22]. In addition, in the calculation of the $C^{2}$-estimate of $\phi$ and the gradient estimate of $F$, we need to modify Chen-Cheng’s method in [CC21b] (in particular, we need to choose new test functions and integral weights in order to deal with the third-order derivative terms (of $\phi$)).
@@
In Section 4, we prove Theorem 1.2 by using the continuity method. We give a sketched proof of the openness part in Subsection 4.A, where we follow the argument in [Che18], [Has19]. We prove the closedness part in Subsection 4.B, where we follow the strategy of [CC21b].
@@ Acknowledgement
The authors would like to thank Abdellah Lahdili for sharing his idea on showing the openness near the starting point of the continuity path around 2020. The authors would like to thank Eleonora Di Nezza, Simon Jubert, Abdellah Lahdili for pointing out an error (which causes a problem in $W^{2,p}$ estimate if we don’t assume the log-concavity of $\mathrm{v}$) in the first version of our draft. The authors would like to thank Jianchun Chu, Lifan Guan, Chang Li for their helpful discussions. The authors would like to thank Prof. Gang Tian, Akito Futaki, Chi Li, Eiji Inoue, Kewei Zhang for their interest and helpful comments. J. Han is supported by National Key R&D Program of China 2023YFA1009900, NSFC-20701003A012401.
@@ 2.A. Weighted Monge-Ampère measures, space of potentials, and functionals
The variational approach has successful applications in the resolution of Kähler-Einstein problem, see [Tia97, BBGZ13, BBE+19, BBJ21]. In this subsection, we will define the notions needed in the variational approach. We will follow the lines in [BN14, HL23, AJL23].
@@
[…]
In [BN14], the authors define the weighted Monge-Ampère measure by using Kiselman’s partial Legendre transform and approximation by series of step functions. Then in [HL23], the authors present an alternative definition by using a more direct fibration construction, which is more adapted in defining the non-Archimedean functionals.
Proposition 2.1 ([BN14, HL23]).
@@
The following weighted functionals and potential spaces are defined in [BN14, HL23, Lah19, AJL23].
@@
${\mathbb{E}}_{\mathrm{v}}^{\eta}(\phi,\phi_{0})=\frac{1}{\mathbb{V}_{\mathrm{v}}(X)}\int_{0}^{1}\int_{X}(\phi-\phi_{0})\Big{(}\mathrm{v}(m_{\phi_{t}})\eta\wedge\frac{\omega_{\phi_{t}}^{n-1}}{(n-1)!}+\langle\mathrm{d}\mathrm{v}(m_{\phi_{t}}),m_{\eta}\rangle_{\phi_{t}}\frac{\omega_{\phi_{t}}^{n}}{n!}\Big{)}\mathrm{d}t,$ (2.3)
where $\eta$ is a smooth $(1,1)$-form, which can be locally expressed as $\eta={\rm dd^{c}}\psi$. And $m_{\eta}(\xi):=-J\xi(\psi)$, which is globally well-defined. And $\langle d\mathrm{v}(m_{\phi},m_{\eta}\rangle_{\phi}=-\sum_{\alpha}\mathrm{v}_{,\alpha}J\xi^{\alpha}(\psi)$.
@@
$\frac{1}{C}\big{(}{\mathbb{I}}(\phi)-{\mathbb{J}}(\phi)\big{)}\leq{\mathbb{I}}_{\mathrm{v}}(\phi)-{\mathbb{J}}_{\mathrm{v}}(\phi)\leq C\big{(}{\mathbb{I}}(\phi)-{\mathbb{J}}(\phi)\big{)}.$
@@
In [Che00, Dar15, BN14, BB17, BDL17, Lah19, HL23, Lah23], the authors show that the space ${\mathcal{E}}^{1}_{K}(X,\omega)$ is granted with a Finsler structure. The functionals have controlled behaviours with respect to such Finsler structure.
@@
All the definitions and propositions of this subsection can be generalized to the setting of projective klt varieties. See for example [HL23, Section 2].
@@ 2.B. Weighted-cscK equations
In this subsection, following the idea in [CC21a], we re-write weighted-cscK equation (1.1) to a coupled PDE system, in which the first one is a $g$-Monge-Ampère equation (as studied in [BN14], [HL23]) and the second one is a linear elliptic equation. Compared with the cscK equation studied in [CC21a], one main difference is that, there are first order derivative terms appearing in the second equation. In the following of this subsection, we will collect some computations that are frequently used in latter sections.
@@
$F=\log\left(\frac{\mathrm{v}(m_{\phi})\omega_{\phi}^{n}}{\omega^{n}}\right)\quad i.e.,\mathrm{v}(m_{\phi})\omega_{\phi}^{n}=e^{F}\omega^{n}.$
@@
$\Delta_{\phi}\log\mathrm{v}(m_{\phi})=\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m^{\xi^{\alpha}}_{\phi})+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$ (2.10)
@@
$=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}+\mathrm{v}(m_{\phi})^{-1}\Big{(}\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\mathrm{Ric}(\omega)}\rangle+\mathrm{v}(m_{\phi})\mathrm{tr}_{\phi}\mathrm{Ric}(\omega)\Big{)}$
$=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})}$
@@
$S_{\mathrm{v}}(\phi)=\mathrm{v}(m_{\phi})S(\phi)-2\mathrm{v}_{,\alpha}\Delta_{\phi}(m_{\phi}^{\xi^{\alpha}})-\mathrm{v}_{,\alpha\beta}(m_{\phi})\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}.$
and $m_{\mathrm{Ric}(\omega)}:=\Delta_{\omega}(m_{\omega})$.
@@
$-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\beta},\xi^{\alpha}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m_{\phi}^{\xi^{\alpha}}-m_{\omega}^{\xi^{\alpha}})$
@@
$=$ $-\left((\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}\right)+\mathrm{tr}_{\phi}\eta_{t}$
@@
$f_{t}^{\prime}:=(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}}).$
In addition, by (2.10) and (2.14)
@@
$f_{t}=(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}.$
@@
$\qquad\qquad\qquad\qquad\qquad\quad-\frac{\mathrm{tr}_{\mathrm{v},\phi}({\rm dd^{c}}f_{*})}{\mathrm{v}(m_{\phi})}-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})},$ (2.20b)

### v3

@@ 1. Introduction
In this paper, we apply the variational approach and continuity method to study the weighted-cscK (extremal) metrics, the study of which is initiated in papers [Lah19, Lah23, AJL23]. The weighted-cscK (extremal) metric is a natural generalization of the cscK (extremal) metric. The (generalized) Yau–Tian–Donaldson conjecture predicts an equivalence between the existence of cscK metric(s) and the $\mathbb{G}$}-uniform K-stability condition. By the work of many authors (see an incomplete list: [Tian97, DR17, BDL20, CC21a, CC21b, Li22a, BJ23]), much progress has been made, and it is known that the existence of cscK metric(s) is equivalent to the $\mathbb{G}$}-coercivity of the Mabuchi functional, and can be implied by the $\mathbb{G}$}-uniformly K-stability for models.
@@
$\mathrm{tr}_{\mathrm{v},\phi}(\eta):=\mathrm{v}(m_{\phi})\mathrm{tr}_{\phi}\eta+\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\eta}\rangle,$
where the definition of $\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\eta}\rangle$ can be found in Definition 2.2.
@@
To show (2) implies (1), we consider the following weighted version of the continuity path of Chen [Ch18] for weighted extremal metrics,
@@
where $\theta$ is a $K$-invariant smooth positive $(1,1)$-form in the same class of $\omega$ and $\underline{\theta}$ is defined by (4.2), which is a constant only depending on $\omega]$ by [Lah19, Lemma 2].
@@
The idea of the continuity path (1.3) can be seen from the variational picture. Since $\mathrm{Ric}(\omega)$ may not be positive, replacing a small part of the term ${\mathbb{E}}_{\mathrm{v}}^{\mathrm{Ric}}$ in the weighted Mabuchi functional by ${\mathbb{E}}_{\mathrm{v}}^{\theta}$ for some positive (1,1)-form $\theta$ will make the weighted Mabuchi functional more coercive. Intuitively, this makes (1.3) always solvable for $t\in[0,1)$.
@@
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})}.$ (1.4b)
@@
We prove a uniform $C^{0}$-estimate by modifying Guo-Phong’s argument in [guo2022]. In addition, in the calculation of the $C^{2}$-estimate of $\phi$ and the gradient estimate of $F$, we need to modify Chen-Cheng’s method in [CC21b] (in particular, we need to choose new test functions and integral weights in order to deal with the third-order derivative terms (of $\phi$)).
@@
In Section 4, we prove Theorem 1.2 by using the continuity method. We give a sketched proof of the openness part in Subsection 4.A, where we follow the argument in [Ch18], [Has19]. We prove the closedness part in Subsection 4.B, where we follow the strategy of [CC21b].
@@ Acknowledgement
The authors would like to thank Abdellah Lahdili for sharing his idea on showing the openness near the starting point of the continuity path around 2020. The authors would like to thank Eleonora Di Nezza, Simon Jubert, Abdellah Lahdili for pointing out an error (which causes a problem in $W^{2,p}$ estimate if we don’t assume the log-concavity of $\mathrm{v}$) in the first version of our draft. The authors would like to thank Jianchun Chu, Lifan Guan, Chang Li for their helpful discussions. The authors would like to thank Prof. Gang Tian, Akito Futaki, Chi Li, Eiji Inoue, Kewei Zhang for their interest and helpful comments. The authors would like to thank the anonymous referees for many helpful comments. J. Han is supported by National Key R&D Program of China 2023YFA1009900, NSFC No.12301059, XHD23A0101.
@@ 2.A. Weighted Monge-Ampère measures, space of potentials, and functionals
The variational approach has successful applications in the resolution of Kähler-Einstein problem, see [Tian97, BBGZ13, BBEGZ19, BBJ21]. In this subsection, we will define the notions needed in the variational approach. We will follow the lines in [BWN14, HL20, AJL23].
@@
$\mathrm{Psh}_{K}(X,\omega)=\{\phi\in\mathrm{Psh}(X,\omega):\phi\text{ is invariant under the }K\text{-action}\}.$ (2.1)
In [BWN14], the authors define the weighted Monge-Ampère measure by using Kiselman’s partial Legendre transform and approximation by series of step functions. Then in [HL20], the authors present an alternative definition by using a more direct fibration construction, which is more adapted in defining the non-Archimedean functionals.
Proposition 2.1 ([BWN14, HL20]).
@@
It should be noted that such weighted Monge–Ampère measure is understood in the sense of non-pluripolar product. The following weighted functionals and potential spaces are defined in [BWN14, HL20, Lah19, AJL23].
@@
${\mathbb{E}}_{\mathrm{v}}^{\eta}(\phi,\phi_{0})=\frac{1}{\mathbb{V}_{\mathrm{v}}(X)}\int_{0}^{1}\int_{X}(\phi-\phi_{0})\Big(\mathrm{v}(m_{\phi_{t}})\eta\wedge\frac{\omega_{\phi_{t}}^{n-1}}{(n-1)!}+\langle\mathrm{d}\mathrm{v}(m_{\phi_{t}}),m_{\eta}\rangle\frac{\omega_{\phi_{t}}^{n}}{n!}\Big)\mathrm{d}t,$ (2.3)
where $\eta$ is a smooth $(1,1)$-form, which can be locally expressed as $\eta={\rm dd^{c}}\psi$. And $m_{\eta}(\xi):=-J\xi(\psi)$, which is globally well-defined. And $\langle\mathrm{d}\mathrm{v}(m_{\phi}),\cdot\rangle=\mathrm{d}\mathrm{v}(m_{\phi})^{\sharp}=\sum_{\alpha}\mathrm{v}_{,\alpha}\xi^{\alpha}$ as a $\mathfrak{t}$-valued function on $X$, where $(\cdot)^{\sharp}$ is with respect to the Kähler form. Therefore, $\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\eta}\rangle=\sum_{\alpha}\mathrm{v}_{,\alpha}\cdot\xi^{\alpha}(m_{\eta})=-\sum_{\alpha}\mathrm{v}_{,\alpha}J\xi^{\alpha}(\psi)$.
@@
$\frac{1}{C}\big({\mathbb{I}}(\phi)-{\mathbb{J}}(\phi)\big)\leq{\mathbb{I}}_{\mathrm{v}}(\phi)-{\mathbb{J}}_{\mathrm{v}}(\phi)\leq C\big({\mathbb{I}}(\phi)-{\mathbb{J}}(\phi)\big),$
where $\mathbb{I}$}, $\mathbb{J}$} are the classical $I$-functional and $J$-functional:
${\mathbb{I}}(\phi):=\frac{1}{\mathbb{V}(X)}\int_{X}\phi\omega^{n}-\frac{1}{\mathbb{V}(X)}\int_{X}\phi\omega_{\phi}^{n}$
and
${\mathbb{J}}(\phi):=\frac{1}{\mathbb{V}(X)}\int_{X}\phi\omega^{n}-E(\phi)=\frac{1}{\mathbb{V}(X)}\int_{X}\phi\omega^{n}-\frac{1}{(n+1)\mathbb{V}(X)}\sum_{j=0}^{n}\int_{X}\phi\omega_{\phi}^{j}\wedge\omega^{n-j}.$
@@
In [Chen00, Dar15, BWN14, BB17, BDL17, Lah19, HL20, Lah23], the authors show that the space ${\mathcal{E}}^{1}_{K}(X,\omega)$ is granted with a Finsler structure. The functionals have controlled behaviours with respect to such Finsler structure.
@@
All the definitions and propositions of this subsection can be generalized to the setting of projective klt varieties. See for example [HL20, Section 2].
@@ 2.B. Weighted-cscK equations
In this subsection, following the idea in [CC21a], we re-write weighted-cscK equation (1.1) to a coupled PDE system, in which the first one is a $g$-Monge-Ampère equation (as studied in [BWN14], [HL20]) and the second one is a linear elliptic equation. Compared with the cscK equation studied in [CC21a], one main difference is that, there are first order derivative terms appearing in the second equation. In the following of this subsection, we will collect some computations that are frequently used in latter sections.
@@
$F=\log\left(\frac{\mathrm{v}(m_{\phi})\omega_{\phi}^{n}}{\omega^{n}}\right)\qquad i.e.,\mathrm{v}(m_{\phi})\omega_{\phi}^{n}=e^{F}\omega^{n}.$
@@
$\Delta_{\phi}\log\mathrm{v}(m_{\phi})=$ $g_{\phi}^{i\overline{j}}\partial_{i}\partial_{\overline{j}}(\log\mathrm{v}(m_{\phi}))=g_{\phi}^{i\overline{j}}\partial_{i}\frac{\partial_{\overline{j}}\mathrm{v}(m_{\phi})}{\mathrm{v}(\phi)}=g_{\phi}^{i\overline{j}}\partial_{i}\frac{\mathrm{v}_{,\alpha}(m_{\phi})\partial_{\overline{j}}m_{\phi}^{\xi^{\alpha}}}{\mathrm{v}(\phi)}$
$=$ $g_{\phi}^{i\overline{j}}\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\partial_{i}\partial_{\overline{j}}m_{\phi}^{\xi^{\alpha}}+g_{\phi}^{i\overline{j}}\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})\partial_{i}m_{\phi}^{\xi^{\beta}}\partial_{\overline{j}}m_{\phi}^{\xi^{\alpha}}}{\mathrm{v}(\phi)}-g_{\phi}^{i\overline{j}}\frac{\mathrm{v}_{,\beta}(m_{\phi})\partial_{i}m_{\phi}^{\xi^{\alpha}}\mathrm{v}_{,\alpha}(m_{\phi})\partial_{\overline{j}}m_{\phi}^{\xi^{\alpha}}}{\mathrm{v}(\phi)^{2}}$
$=$ $\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\phi}(m^{\xi^{\alpha}}_{\phi})+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$ (2.11)
@@
$=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}+\mathrm{v}(m_{\phi})^{-1}\Big(\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\mathrm{Ric}(\omega)}\rangle+\mathrm{v}(m_{\phi})\mathrm{tr}_{\phi}\mathrm{Ric}(\omega)\Big)$
$=$ $-\frac{S_{\mathrm{v}}(\phi)}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})},$
@@
$S_{\mathrm{v}}(\phi)=\mathrm{v}(m_{\phi})S(\phi)-2\mathrm{v}_{,\alpha}\Delta_{\phi}(m_{\phi}^{\xi^{\alpha}})-\mathrm{v}_{,\alpha\beta}(m_{\phi})\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi},$
$m_{\mathrm{Ric}(\omega)}=\Delta_{\omega}(m_{\omega})$ and
$-\mathrm{v}_{,\alpha}(m_{\phi})\Delta_{\omega}(m^{\xi^{\alpha}}_{\omega})=\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\mathrm{Ric}(\omega)}\rangle$
by [Lah19, Lemma 5]. Indeed, locally we can write $\omega={\rm dd^{c}}\psi$, $\mathrm{Ric}(\omega)=-{\rm dd^{c}}\log\det(\psi_{i\overline{j}})$, and Ricci potential is $\log\det(\psi_{i\overline{j}})$. Then $\Delta_{\omega}(m^{\xi^{\alpha}}_{\omega})=g^{i\overline{j}}\partial_{i\overline{j}}(-J\xi^{\alpha}\psi)=J\xi^{\alpha}(-\log\det(\psi_{i\overline{j}}))=m_{\mathrm{Ric}(\omega)}$.
@@
$-\frac{\mathrm{v}_{,\alpha}(m_{\phi})\mathrm{v}_{,\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})^{2}}\langle\xi^{\beta},\xi^{\alpha}\rangle_{\phi}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}(\Delta_{\phi}m_{\phi}^{\xi^{\alpha}}-\Delta_{\omega}m_{\omega}^{\xi^{\alpha}})$
@@
$=$ $-\left((\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}\right)+\mathrm{tr}_{\phi}\eta_{t}$
@@
$f_{t}^{\prime}:=(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}+2\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}}).$
In addition, by (2.B) and (2.15)
@@
$f_{t}=(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}-(1-\frac{1}{t})\frac{\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}.$
@@
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\qquad-\frac{\mathrm{tr}_{\mathrm{v},\phi}({\rm dd^{c}}f_{*})}{\mathrm{v}(m_{\phi})}-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi})}{\mathrm{v}(m_{\phi})}+\frac{\mathrm{tr}_{\mathrm{v},\phi}\mathrm{Ric}(\omega)}{\mathrm{v}(m_{\phi})},$ (2.21b)